# Sim-to-real supervised domain adaptation for radioisotope identification

### v3

@@ 4.3 Hyperparameter Search
Prior to training, we performed an in-depth Bayesian hyperparameter search using the Optuna software package [38]. We determined architecture hyperparameters using the source domain dataset in the sim-to-sim scenario with validation loss as our search criterion. To ensure a direct comparison, these architecture hyperparameters were used for all models in this study, while training hyperparameters (learning rate, batch size, dropout, weight decay) were calculated separately for each domain adaptation scenario. Since we might expect the optimal training hyperparameters to vary with dataset size, we performed a separate hyperparameter search for every data subset size. The fine-tuning hyperparameters were determined separately from the target-only hyperparameters, and the fine-tuning runs considered which (if any) layers to freeze as part of the hyperparameter search. Similar to the results of Ref. [39], we sometimes found that freezing the output layer(s) while fine-tuning only the earlier layers yielded better performance. We provide a further discussion of the hyperparameter tuning process in Section A.2.
@@ 4.4 Uncertainty Estimation
We employed an ensemble method to estimate the uncertainty in our model performance. Each model was trained with 10 different random initial configurations, and the variations in performance were used to quantify model uncertainty. For the source-only model class, this simply meant training 10 models for each architecture using different random weight initializations and seed configurations. For the target-only and domain-adapted model classes, we additionally sampled a random subset of target-domain data during each random trial. These trials were paired by ensuring that the random data subset was identical between target-only and domain-adapted models with the same seed. Similarly, each fine-tuning trial was initialized from the source-only pretraining checkpoint corresponding to its seed, so the measured variance reflects both pretraining and fine-tuning.
@@ 5.1 Predictive Performance (Classification Metrics)
After training three classes of models (source-only, target-only, domain-adapted) for each of the three domain adaptation scenarios (sim-to-sim HPGe, sim-to-real LaBr3, sim-to-real NaI(Tl)) with different architectural backbones (MLP, CNN, TBNN (Li), TBNN (ours)) for a range of different dataset sizes, we recorded the performance of each model on a held-out target-domain testing dataset. The choice of a best metric to be used for spectroscopic analysis has no consensus in the literature, although the detection limit and quantification uncertainty are preferred by spectroscopists for the detection and quantification of individual isotopes [4, 14, 40]. Without selecting details about the amount of background present and the isotopes of interest, these metrics are not computable, and traditional machine learning metrics characterizing the classifiers themselves must be used111Unfortunately, some evidence in [14] shows that accuracy may not be monotonic with detection limit or quantification accuracy, so caution should be used when extending classifier metrics to real-world performance.. For the sim-to-real scenarios, we used testing accuracy as our evaluation metric. For the sim-to-sim scenario, accuracy is no longer a well-defined metric since we are performing multi-label proportion estimation; instead, we computed the testing absolute proportion error (APE) score as our evaluation metric. The APE score (Eq. 1) provides a convenient 0-to-1 scoring metric to compare how precisely different models predict the fractional composition of each mixed source. We provide a more detailed discussion of the APE score as an evaluation metric in Section A.1.
@@ 5.1.1 Sim-to-sim comparison
We summarize the results of the sim-to-sim scenario in Fig. 3, where we compare the geant APE score across different model classes for each architecture. Overall, we observe that the APE scores of the domain-adapted models very often match or exceed those of the source-only and target-only models. In particular, using a fine-tuning dataset of 1024 spectra, the TBNN (ours) model achieves a testing APE score of , outperforming both the source-only model (APE score = ) and the target-only model (APE score = ) on the same dataset size. We quantify these results in Table 2(a) where we explicitly record the testing APE scores for models trained with 0 versus 64 versus 1024 target domain training spectra.
@@ 5.1.2 Sim-to-real comparison
In the sim-to-real scenarios (Figs. 6(c) and 6(b)), we do not observe a statistically significant difference between model architectures for most dataset sizes. As indicated in Tables 7(b) and 7(c), there is not a strong, consistent trend supporting one model architecture over the others. We hypothesize that the relative simplicity of the sim-to-real classification problem offers minimal advantages to using more complex architectures, and thus recommend that future work perform similar architectural comparisons on more challenging experimental classification scenarios.
@@ 5.2.1 Model Diagnostics
We present our results in Table 3 using a TBNN (ours) architecture for the sim-to-real LaBr3 adaptation scenario. In addition to standard classification accuracy (acc), we include negative log-likelihood (nll), which more heavily penalizes overconfidence [41]. The Brier score (brier) measures the mean-squared error between the true one-hot encoded label and predicted probability vectors [42], while the expected calibration error (ece) bins predictions by confidence to directly measure the confidence-accuracy gap [43]. Prediction margin evaluates the mean (margin_mean) and 10th percentile (margin_p10) of the margin for each spectrum, defined as the difference in predicted probabilities between the top two classes [44]. Mean predictive entropy (entropy_mean) measures how confident or uncertain the model is in its predictions by calculating the average Shannon entropy of the predicted probability vectors [45]. Mean Jacobian norm (jacobian_norm_mean) quantifies a model’s sensitivity to its input by averaging the squared -norm of the gradient of the loss with respect to the input across all samples [46]. Lastly, we considered four boundary complexity and smoothness metrics by first building a -NN graph using a Euclidean distance metric and on the input -normalized spectra of the testing dataset. We considered a range of distance metrics (Euclidean, cosine, correlation) and different values for (5-30), finding minimal changes in results. We then considered the post-softmax probability vectors for each edge and computed the following metrics [47, 48]:
@@
Metric Source-only Target-only 64 Domain-adapted 64 -value (best vs. 2nd) Dir. Type acc Accuracy nll [41] Likelihood / Loss brier [42] Calibration ece [43] Calibration margin_mean [44] Separation / Margin margin_p10 [44] Separation / Margin entropy_mean [45] Uncertainty jacobian_norm_mean [46] Sensitivity knn_tv_hard [47, 48] Boundary Complexity knn_prob_l2 [47, 48] Smoothness knn_conf_absdiff [47, 48] Smoothness knn_margin_absdiff [47, 48] Smoothness
@@ 5.2.2 Model Interpretability
We also present evidence that beyond metric scores, domain-adapted models learn more human-interpretable features compared to target-only models. In Explainable AI literature, various techniques have been proposed for quantifying input feature relevance, such as gradient-based [49, 50] and Attention Rollout [51] methods. In Ref. [52], the authors considered various explanation methods in the context of gamma spectroscopy, and concluded that SHAP offered especially accurate explanations of radioisotope predictions compared to other methods. In short, SHAP evaluates an input feature’s importance by removing the feature and assessing the resulting effect on the model’s output. A more mathematically thorough description of the KernelSHAP algorithm can be found in Refs. [52, 53].
@@
For this analysis, we chose a target-only and domain-adapted TBNN (ours) model from the sim-to-real LaBr3 adaptation scenario, each trained with the same 64 experimental spectra. Fig. 2 shows a 99Mo spectrum overlaid with a color plot of calculated SHAP explanation values for each region. Large, positive SHAP values (highlighted in red) indicate that a feature positively correlates with the model’s class prediction. The SHAP explanation reveals that the target-only model relies exclusively on the 140.5 keV peak from 99mTc when computing its classification, ignoring the 739.5 keV peak characteristic of 99Mo. As a result, the target-only model frequently confuses 99mTc and 99Mo due to reliance on the shared daughter line, resulting in an overall false alarm rate of 2.0% for 99Mo. Conversely, the domain-adapted model correctly identifies both the 140.5 keV and 739.5 keV peaks, enabling unambiguous discrimination between both isotopes.
@@
Our analysis further reveals that target-only models will often overfit to the low energy region of the spectrum. As seen in Fig. 7, when classifying 239Pu, the target-only model identifies the - keV band as the most salient component of the spectrum, indicative of overfitting. Conversely, the domain-adapted model’s SHAP explanation most strongly highlights a region containing the 345, 375, and 413 keV peaks, which are commonly used by spectroscopists to identify 239Pu, especially in the presence of shielding [54]. Overall, we observe that domain-adapted models more consistently produce explanations that align with human-interpretable photopeaks, while target-only models are more likely to focus on low-energy, high-count features. As such, we hypothesize that gadras pretraining confers an interpretability boost, even when the target-only and domain-adapted models have similar per-class accuracy.
@@ A.2 Hyperparameter search
For each architecture (MLP, CNN, TBNN (Li), TBNN (ours)), we conducted a Bayesian hyperparameter search with 150 trials on the sim-to-sim source domain dataset. Each model was trained for 20 minutes on an NVIDIA Tesla V100 GPU with 32 GB of memory, with the number of epochs adjusted dynamically to maintain a consistent overall training time. Underperforming configurations were pruned early, and an early stopping criterion was applied to mitigate overfitting. Hyperparameters were ranked by their validation cross-entropy loss, and the best-performing parameters are shown in Table 4. Source-only baseline models for all scenarios and all architectures were trained for 80 minutes on the same hardware, and the training was repeated 10 times using different random seeds.
@@
(a) Training hyperparameters for the sim-to-sim gadrasgeant adaptation using a HPGe detector. Best run (target-only) Best run (domain-adapted) Architecture Parameter Search Space N=64 N=1024 N=64 N=1024 MLP Learning Rate {1e-6, 1e-3} 9.58e-5 9.01e-4 2.00e-4 3.82e-4 Batch Size {32, 512} 32 512 64 512 Weight Decay {1e-7, 1e-1} 4.08e-4 2.92e-3 3.17e-2 3.69e-5 Dropout {0.0, 0.4} 0.373 0.337 0.397 0.395 Frozen Layers {none, all} N/A N/A None None CNN Learning Rate {1e-6, 1e-3} 1.92e-5 7.16e-4 9.68e-4 8.71e-5 Batch Size {32, 512} 32 512 64 64 Weight Decay {1e-7, 1e-1} 7.76e-2 2.84e-4 9.84e-7 5.36e-7 Dropout {0.0, 0.4} 0.381 0.275 0.303 0.322 Frozen Layers {none, all} N/A N/A Last 3 layers First 2 layers TBNN (Li) Learning Rate {1e-6, 1e-3} 6.73e-4 9.75e-4 6.44e-4 6.53e-4 Batch Size {32, 512} 64 128 32 512 Weight Decay {1e-7, 1e-1} 2.73e-7 2.14e-2 1.71e-5 6.86e-5 Dropout {0.0, 0.4} 0.149 0.067 0.027 0.021 Frozen Layers {none, all} N/A N/A None None TBNN (ours) Learning Rate {1e-6, 1e-3} 3.37e-4 7.67e-4 1.82e-4 5.77e-5 Batch Size {32, 512} 32 32 64 512 Weight Decay {1e-7, 1e-1} 1.31e-2 4.43e-5 5.48e-4 1.36e-6 Dropout {0.0, 0.4} 0.051 0.159 0.175 0.186 Frozen Layers {none, all} N/A N/A None Last 5 layers (b) Training hyperparameters for the sim-to-real gadrasexperiment adaptation using a LaBr3 detector. Best run (target-only) Best run (domain-adapted) Architecture Parameter Search Space N=64 N=1024 N=64 N=1024 MLP Learning Rate {1e-6, 1e-3} 1.19e-6 1.44e-6 8.10e-6 2.78e-4 Batch Size {32, 512} 64 32 64 256 Weight Decay {1e-7, 1e-1} 1.45e-6 8.26e-2 4.24e-7 3.80e-5 Dropout {0.0, 0.4} 0.021 0.098 0.000312 0.344 Frozen Layers {none, all} N/A N/A Last 2 layers Last layer CNN Learning Rate {1e-6, 1e-3} 2.69e-6 5.06e-5 1.13e-4 4.49e-4 Batch Size {32, 512} 64 32 32 256 Weight Decay {1e-7, 1e-1} 9.62e-5 3.58e-2 1.06e-5 2.33e-5 Dropout {0.0, 0.4} 0.000239 0.0349 0.00476 0.389 Frozen Layers {none, all} N/A N/A First 3 layers None TBNN (Li) Learning Rate {1e-6, 1e-3} 2.65e-5 8.79e-4 5.07e-4 8.85e-4 Batch Size {32, 512} 32 128 32 256 Weight Decay {1e-7, 1e-1} 2.62e-4 6.98e-5 1.10e-2 1.30e-4 Dropout {0.0, 0.4} 0.0117 0.0125 0.156 0.0125 Frozen Layers {none, all} N/A N/A Last layer Last 8 layers TBNN (ours) Learning Rate {1e-6, 1e-3} 6.95e-6 2.25e-4 7.02e-4 2.37e-4 Batch Size {32, 512} 32 32 32 128 Weight Decay {1e-7, 1e-1} 1.63e-4 2.76e-5 4.54e-7 1.49e-4 Dropout {0.0, 0.4} 0.000792 0.0347 0.0476 0.370 Frozen Layers {none, all} N/A N/A Last 19 layers Last layer (c) Training hyperparameters for the sim-to-real gadrasexperiment adaptation using a NaI(Tl) detector. Best run (target-only) Best run (domain-adapted) Architecture Parameter Search Space N=64 N=1024 N=64 N=1024 MLP Learning Rate {1e-6, 1e-3} 1.01e-6 7.79e-6 2.37e-4 4.84e-4 Batch Size {32, 512} 32 32 64 32 Weight Decay {1e-7, 1e-1} 4.05e-7 1.78e-2 1.49e-4 4.03e-4 Dropout {0.0, 0.4} 0.012 0.019 0.370 0.141 Frozen Layers {none, all} N/A N/A None Last layer CNN Learning Rate {1e-6, 1e-3} 2.14e-6 1.66e-4 7.43e-6 2.27e-4 Batch Size {32, 512} 64 256 64 32 Weight Decay {1e-7, 1e-1} 2.32e-3 4.70e-3 4.61e-5 4.51e-5 Dropout {0.0, 0.4} 0.00335 0.00206 0.0120 0.00117 Frozen Layers {none, all} N/A N/A Last layer None TBNN (Li) Learning Rate {1e-6, 1e-3} 6.39e-4 3.75e-4 1.84e-4 7.74e-4 Batch Size {32, 512} 64 64 64 32 Weight Decay {1e-7, 1e-1} 1.30e-4 4.65e-4 2.33e-4 7.06e-4 Dropout {0.0, 0.4} 0.00170 0.0298 0.00506 0.0383 Frozen Layers {none, all} N/A N/A First 5 layers First 13 layers TBNN (ours) Learning Rate {1e-6, 1e-3} 9.29e-5 2.35e-4 7.31e-4 2.53e-4 Batch Size {32, 512} 64 32 32 64 Weight Decay {1e-7, 1e-1} 2.13e-2 6.41e-7 4.98e-2 6.62e-5 Dropout {0.0, 0.4} 0.00192 0.0164 0.184 0.350 Frozen Layers {none, all} N/A N/A First 3 layers None

### v4

@@ 4.3 Hyperparameter Search
Prior to training, we performed an in-depth Bayesian hyperparameter search using the Optuna software package [optuna2019]. We determined architecture hyperparameters using the source domain dataset in the sim-to-sim scenario with validation loss as our search criterion. To ensure a direct comparison, these architecture hyperparameters were used for all models in this study, while training hyperparameters (learning rate, batch size, dropout, weight decay) were calculated separately for each domain adaptation scenario. Since we might expect the optimal training hyperparameters to vary with dataset size, we performed a separate hyperparameter search for every data subset size. The fine-tuning hyperparameters were determined separately from the target-only hyperparameters, and the fine-tuning runs considered which (if any) layers to freeze as part of the hyperparameter search. Similar to the results of Ref. [shirokikh2020first], we sometimes found that freezing the output layer(s) while fine-tuning only the earlier layers yielded better performance. We provide a further discussion of the hyperparameter tuning process in Section A.2.
@@ 4.4 Uncertainty Estimation
We employed an ensemble method to estimate the uncertainty in our model performance. Each model was trained with 10 different random initial configurations, and the variations in performance were used to quantify model uncertainty. For the source-only model class, this simply meant training 10 models for each architecture using different random weight initializations and seed configurations. For the target-only and domain-adapted model classes, we additionally sampled a random subset of target-domain data during each random trial. These trials were paired by ensuring that the random data subset was identical between target-only and domain-adapted models with the same seed. Similarly, each fine-tuning trial was initialized from the source-only pretraining checkpoint corresponding to its seed, so the measured variance reflects both pretraining and fine-tuning. Our reported uncertainty estimate reflects the empirical sample standard deviation across these 10 trials along with an added term that accounts for the sampling variability arising from a finite test set size.
@@ 5.1 Predictive Performance (Classification Metrics)
After training three classes of models (source-only, target-only, domain-adapted) for each of the three domain adaptation scenarios (sim-to-sim HPGe, sim-to-real LaBr3, sim-to-real NaI(Tl)) with different architectural backbones (MLP, CNN, TBNN (Li), TBNN (ours)) for a range of different dataset sizes, we recorded the performance of each model on a held-out target-domain testing dataset. The choice of a best metric to be used for spectroscopic analysis has no consensus in the literature, although the detection limit and quantification uncertainty are preferred by spectroscopists for the detection and quantification of individual isotopes [Hagen2022, fernandezImprovementNuclideDetection2025, Hagen2021]. Without selecting details about the amount of background present and the isotopes of interest, these metrics are not computable, and traditional machine learning metrics characterizing the classifiers themselves must be used111Unfortunately, some evidence in [fernandezImprovementNuclideDetection2025] shows that accuracy may not be monotonic with detection limit or quantification accuracy, so caution should be used when extending classifier metrics to real-world performance.. For the sim-to-real scenarios, we used testing accuracy as our evaluation metric. For the sim-to-sim scenario, accuracy is no longer a well-defined metric since we are performing multi-label proportion estimation; instead, we computed the testing absolute proportion error (APE) score as our evaluation metric. The APE score (Eq. 1) provides a convenient 0-to-1 scoring metric to compare how precisely different models predict the fractional composition of each mixed source. We provide a more detailed discussion of the APE score as an evaluation metric in Section A.1.
@@ 5.1.1 Sim-to-sim comparison
We summarize the results of the sim-to-sim scenario in Fig. 3, where we compare the Geant4 APE score across different model classes for each architecture. Overall, we observe that the APE scores of the domain-adapted models almost always match or exceed those of the source-only and target-only models. In particular, using a fine-tuning dataset of 1024 spectra, the TBNN (ours) model achieves a testing APE score of , outperforming both the source-only model (APE score = ) and the target-only model (APE score = ) on the same dataset size. We quantify these results in Table 2(a) where we explicitly record the testing APE scores for models trained with 0 versus 64 versus 1024 target domain training spectra.
@@ 5.1.2 Sim-to-real comparison
In the sim-to-real scenarios (Figs. 6(c) and 6(b)), we still observe a statistically significant performance gain from transformers at some dataset sizes. However, the contrast is far less widespread, as most dataset sizes show no statistically significant difference between model architectures. This result suggests that, for this particular sim-to-real task, the classification boundary is not sufficiently complex for more expressive architectures (TBNNs) to provide a consistent advantage over classical architectures (MLPs, CNNs). We quantify these results in Tables 7(b) and 7(c), where we find that transformers retain a slight, but statistically significant, edge over classical architectures for some, but not all, dataset sizes. Overall, we recommend that future work repeat such architectural comparisons on more challenging experimental classification scenarios.
@@ 5.2.1 Model Diagnostics
We present our results in Table 3 using a TBNN (ours) architecture for the sim-to-real LaBr3 adaptation scenario. In addition to standard classification accuracy (acc), we include negative log-likelihood (nll), which more heavily penalizes overconfidence [murphy2012machine]. The Brier score (brier) measures the mean-squared error between the true one-hot encoded label and predicted probability vectors [glenn1950verification], while the expected calibration error (ece) bins predictions by confidence to directly measure the confidence-accuracy gap [guo2017calibration]. Prediction margin evaluates the mean (margin_mean) and 10th percentile (margin_p10) of the margin for each spectrum, defined as the difference in predicted probabilities between the top two classes [bartlett1998boosting]. Mean predictive entropy (entropy_mean) measures how confident or uncertain the model is in its predictions by calculating the average Shannon entropy of the predicted probability vectors [shannon1948mathematical]. Mean Jacobian norm (jacobian_norm_mean) quantifies a model’s sensitivity to its input by averaging the squared -norm of the gradient of the loss with respect to the input across all samples [goodfellow2014explaining]. Lastly, we considered four boundary complexity and smoothness metrics by first building a -nearest neighbors graph using a Euclidean distance metric and on the input -normalized spectra of the testing dataset. We considered a range of distance metrics (Euclidean, cosine, correlation) and different values for (5-30), finding minimal changes in results. We then considered the post-softmax probability vectors for each edge and computed the following metrics [belkin2006manifold, zhu2003semi]:
@@
Metric Source-only Target-only 64 Domain-adapted 64 -value (best vs. 2nd) Dir. Type acc Accuracy nll [murphy2012machine] Likelihood / Loss brier [glenn1950verification] Calibration ece [guo2017calibration] Calibration margin_mean [bartlett1998boosting] Separation / Margin margin_p10 [bartlett1998boosting] Separation / Margin entropy_mean [shannon1948mathematical] Uncertainty jacobian_norm_mean [goodfellow2014explaining] Sensitivity knn_tv_hard [belkin2006manifold, zhu2003semi] Boundary Complexity knn_prob_l2 [belkin2006manifold, zhu2003semi] Smoothness knn_conf_absdiff [belkin2006manifold, zhu2003semi] Smoothness knn_margin_absdiff [belkin2006manifold, zhu2003semi] Smoothness
@@ 5.2.2 Model Interpretability
We also present evidence that beyond metric scores, domain-adapted models learn more human-interpretable features compared to target-only models. In Explainable AI literature, various techniques have been proposed for quantifying input feature relevance, such as gradient-based [simonyan2014deepinsideconvolutionalnetworks, smilkov2017smoothgradremovingnoiseadding] and Attention Rollout [abnar2020quantifyingattentionflowtransformers] methods. In Ref. [bandstra2023explaining], the authors considered various explanation methods in the context of gamma spectroscopy, and concluded that SHAP offered especially accurate explanations of radioisotope predictions compared to other methods. In short, SHAP evaluates an input feature’s importance by removing the feature and assessing the resulting effect on the model’s output. A more mathematically thorough description of the KernelSHAP algorithm can be found in Refs. [bandstra2023explaining, lundberg2017unifiedapproachinterpretingmodel].
@@
For this analysis, we chose a target-only and domain-adapted TBNN (ours) model from the sim-to-real LaBr3 adaptation scenario, each trained with the same 64 experimental spectra. Fig. 2 shows a 99Mo spectrum overlaid with a color plot of calculated SHAP explanation values for each region. Large, positive SHAP values (highlighted in red) indicate that a feature positively correlates with the model’s class prediction. The SHAP explanation reveals that the target-only model relies exclusively on the 140.5 keV peak from 99mTc when computing its classification, ignoring the 739.5 keV peak characteristic of 99Mo. As a result, the target-only model frequently confuses 99mTc and 99Mo due to reliance on the shared daughter line, resulting in an overall false alarm rate of 2.0% for 99Mo. Conversely, the domain-adapted model correctly identifies both the 140.5 keV and 739.5 keV peaks, enabling unambiguous discrimination between both isotopes. We report similar findings using other architectures and include the same 99Mo SHAP comparisons for MLP, CNN, and TBNN (Li) architectures in Fig. 7.
@@
Our analysis further reveals that target-only models will often overfit to the low energy region of the spectrum. As seen in Fig. 8, when classifying 239Pu, the target-only model identifies the - keV band as the most salient component of the spectrum, indicative of overfitting. Conversely, the domain-adapted model’s SHAP explanation most strongly highlights a region containing the 345, 375, and 413 keV peaks, which are commonly used by spectroscopists to identify 239Pu, especially in the presence of shielding [reilly1991passive]. Overall, we observe that domain-adapted models more consistently produce explanations that align with human-interpretable photopeaks, while target-only models are more likely to focus on low-energy, high-count features. As such, we hypothesize that GADRAS pretraining confers an interpretability boost, even when the target-only and domain-adapted models have similar per-class accuracy.
We conclude our analysis by examining the behavior of neural network classification models when different isotopes exhibit shared spectral features. Using the domain-adapted TBNN (ours) model, we computed SHAP explanations for a 18F and 22Na spectrum, shown in Fig. 9. The model identifies the 511 keV positron–electron annihilation peak as salient in both spectra; however, the model also highlights the 1274.5 keV line in the 22Na spectrum. This result provides further evidence that the model has learned to utilize secondary features to distinguish between isotopes with shared energy peaks.
@@ A.2 Hyperparameter search
For each architecture (MLP, CNN, TBNN (Li), TBNN (ours)), we conducted a Bayesian hyperparameter search with 150 trials on the sim-to-sim source domain dataset. Each model was trained for 20 minutes on an NVIDIA Tesla V100 GPU with 32 GB of memory, with the number of epochs adjusted dynamically to maintain a consistent overall training time. Underperforming configurations were pruned early, and an early stopping criterion was applied to mitigate overfitting. Hyperparameters were ranked by their validation cross-entropy loss, and the best-performing parameters are shown in Table 4. Source-only baseline models for all scenarios and all architectures were trained for 80 minutes on the same hardware, and the training was repeated 10 times using different random seeds. All trained models were compact ( 2 GB) and achieved single-threaded CPU inference latencies of 2ms, indicating that they could easily be deployed on a commodity system.
@@
(a) Training hyperparameters for the sim-to-sim GADRASGeant4 adaptation using a HPGe detector. Best run (target-only) Best run (domain-adapted) Architecture Parameter Search Space N=64 N=1024 N=64 N=1024 MLP Learning Rate {1e-6, 1e-3} 9.58e-5 9.01e-4 2.00e-4 3.82e-4 Batch Size {32, 512} 32 512 64 512 Weight Decay {1e-7, 1e-1} 4.08e-4 2.92e-3 3.17e-2 3.69e-5 Dropout {0.0, 0.4} 0.373 0.337 0.397 0.395 Frozen Layers {none, all} N/A N/A None None CNN Learning Rate {1e-6, 1e-3} 1.92e-5 7.16e-4 9.68e-4 8.71e-5 Batch Size {32, 512} 32 512 64 64 Weight Decay {1e-7, 1e-1} 7.76e-2 2.84e-4 9.84e-7 5.36e-7 Dropout {0.0, 0.4} 0.381 0.275 0.303 0.322 Frozen Layers {none, all} N/A N/A Last 3 layers First 2 layers TBNN (Li) Learning Rate {1e-6, 1e-3} 6.73e-4 9.75e-4 6.44e-4 6.53e-4 Batch Size {32, 512} 64 128 32 512 Weight Decay {1e-7, 1e-1} 2.73e-7 2.14e-2 1.71e-5 6.86e-5 Dropout {0.0, 0.4} 0.149 0.067 0.027 0.021 Frozen Layers {none, all} N/A N/A None None TBNN (ours) Learning Rate {1e-6, 1e-3} 3.37e-4 7.67e-4 1.82e-4 5.77e-5 Batch Size {32, 512} 32 32 64 512 Weight Decay {1e-7, 1e-1} 1.31e-2 4.43e-5 5.48e-4 1.36e-6 Dropout {0.0, 0.4} 0.051 0.159 0.175 0.186 Frozen Layers {none, all} N/A N/A None Last 5 layers (b) Training hyperparameters for the sim-to-real GADRASexperiment adaptation using a LaBr3 detector. Best run (target-only) Best run (domain-adapted) Architecture Parameter Search Space N=64 N=1024 N=64 N=1024 MLP Learning Rate {1e-6, 1e-3} 1.19e-6 1.44e-6 8.10e-6 2.78e-4 Batch Size {32, 512} 64 32 64 256 Weight Decay {1e-7, 1e-1} 1.45e-6 8.26e-2 4.24e-7 3.80e-5 Dropout {0.0, 0.4} 0.021 0.098 0.000312 0.344 Frozen Layers {none, all} N/A N/A Last 2 layers Last layer CNN Learning Rate {1e-6, 1e-3} 2.69e-6 5.06e-5 1.13e-4 4.49e-4 Batch Size {32, 512} 64 32 32 256 Weight Decay {1e-7, 1e-1} 9.62e-5 3.58e-2 1.06e-5 2.33e-5 Dropout {0.0, 0.4} 0.000239 0.0349 0.00476 0.389 Frozen Layers {none, all} N/A N/A First 3 layers None TBNN (Li) Learning Rate {1e-6, 1e-3} 2.65e-5 8.79e-4 5.07e-4 8.85e-4 Batch Size {32, 512} 32 128 32 256 Weight Decay {1e-7, 1e-1} 2.62e-4 6.98e-5 1.10e-2 1.30e-4 Dropout {0.0, 0.4} 0.0117 0.0125 0.156 0.0125 Frozen Layers {none, all} N/A N/A Last layer Last 8 layers TBNN (ours) Learning Rate {1e-6, 1e-3} 6.95e-6 2.25e-4 7.02e-4 2.37e-4 Batch Size {32, 512} 32 32 32 128 Weight Decay {1e-7, 1e-1} 1.63e-4 2.76e-5 4.54e-7 1.49e-4 Dropout {0.0, 0.4} 0.000792 0.0347 0.0476 0.370 Frozen Layers {none, all} N/A N/A Last 19 layers Last layer (c) Training hyperparameters for the sim-to-real GADRASexperiment adaptation using a NaI(Tl) detector. Best run (target-only) Best run (domain-adapted) Architecture Parameter Search Space N=64 N=1024 N=64 N=1024 MLP Learning Rate {1e-6, 1e-3} 1.01e-6 7.79e-6 2.37e-4 4.84e-4 Batch Size {32, 512} 32 32 64 32 Weight Decay {1e-7, 1e-1} 4.05e-7 1.78e-2 1.49e-4 4.03e-4 Dropout {0.0, 0.4} 0.012 0.019 0.370 0.141 Frozen Layers {none, all} N/A N/A None Last layer CNN Learning Rate {1e-6, 1e-3} 2.14e-6 1.66e-4 7.43e-6 2.27e-4 Batch Size {32, 512} 64 256 64 32 Weight Decay {1e-7, 1e-1} 2.32e-3 4.70e-3 4.61e-5 4.51e-5 Dropout {0.0, 0.4} 0.00335 0.00206 0.0120 0.00117 Frozen Layers {none, all} N/A N/A Last layer None TBNN (Li) Learning Rate {1e-6, 1e-3} 6.39e-4 3.75e-4 1.84e-4 7.74e-4 Batch Size {32, 512} 64 64 64 32 Weight Decay {1e-7, 1e-1} 1.30e-4 4.65e-4 2.33e-4 7.06e-4 Dropout {0.0, 0.4} 0.00170 0.0298 0.00506 0.0383 Frozen Layers {none, all} N/A N/A First 5 layers First 13 layers TBNN (ours) Learning Rate {1e-6, 1e-3} 9.29e-5 2.35e-4 7.31e-4 2.53e-4 Batch Size {32, 512} 64 32 32 64 Weight Decay {1e-7, 1e-1} 2.13e-2 6.41e-7 4.98e-2 6.62e-5 Dropout {0.0, 0.4} 0.00192 0.0164 0.184 0.350 Frozen Layers {none, all} N/A N/A First 3 layers None